# Function spaces for decoupling

### v2

@@ 4.4. Interpolation and duality
$[\mathcal{L}_{W,s_{0}}^{q_{0},p_{0}}(\mathbb{R}^{n}),\mathcal{L}_{W,s_{1}}^{q_{1},p_{1}}(\mathbb{R}^{n})]_{\theta}=\mathcal{L}_{W,s}^{q,p}(\mathbb{R}^{n}),$
@@
It suffices to consider the case where $s_{0}=s_{1}=s=0$. It is a basic fact about interpolation of vector-valued function spaces (see [24, Theorem 2.2.6]) that
$[L^{q_{0}}_{\omega}L^{p_{0}}_{x}L^{2}_{\sigma}(S^{*}_{+}\mathbb{R}^{n}),L^{q_{1}}_{\omega}L^{p_{1}}_{x}L^{2}_{\sigma}(S^{*}_{+}\mathbb{R}^{n})]_{\theta}=L^{q}_{\omega}L^{p}_{x}L^{2}_{\sigma}(S^{*}_{+}\mathbb{R}^{n}).$
Hence one can combine Theorem 3.4 with a result about interpolation of complemented subspaces (see [38, Theorem 1.17.1.1]) to conclude that
$[WVL^{q_{0}}_{\omega}L^{p_{0}}_{x}L^{2}_{\sigma}(S^{*}_{+}\mathbb{R}^{n}),WVL^{q_{1}}_{\omega}L^{p_{1}}_{x}L^{2}_{\sigma}(S^{*}_{+}\mathbb{R}^{n})]_{\theta}=WVL^{q}_{\omega}L^{p}_{x}L^{2}_{\sigma}(S^{*}_{+}\mathbb{R}^{n}).$
@@
$W:\mathcal{L}_{W,0}^{\tilde{q},\tilde{p}}(\mathbb{R}^{n})\to WVL^{\tilde{q}}_{\omega}L^{\tilde{p}}_{x}L^{2}_{\sigma}(S^{*}_{+}\mathbb{R}^{n})$
@@
Conversely, let $l\in(\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n}))^{*}$. Then $l\circ V\in(L^{q}_{\omega}L^{p}_{x}L^{2}_{\sigma}(S^{*}_{+}\mathbb{R}^{n}))^{*}$, by Corollary 4.4. Moreover, by [24, Theorems 1.3.10 and 1.3.21],
@@ 5.1. Boundedness of the Euclidean half-wave group
We first prove that our function spaces are invariant under the action of the Euclidean half-wave propagators. The following theorem contains the first statement of Theorem 1.1.
@@
Let $\rho^{\prime}\in C^{\infty}_{c}(\mathbb{R}^{n})$ be such that $\rho^{\prime}\equiv 1$ on $\mathrm{supp}(\rho)$, where $\rho$ is as in (4.1). It then follows from straightforward kernel estimates (see also [8, Theorem 1.18]) that
@@
for an implicit constant independent of $\omega\in S^{n-1}$ and $t\in\mathbb{R}$. To this end, we will adapt an argument from [15, Section 9] to the setting of anisotropic Hardy spaces.
@@ 5.2. Unboundedness of Fourier integral operators when $p\neq q$
We recall from [19] (see also [27, Corollary 2.4] and Corollary 6.6 below) that $\mathcal{L}_{W,s}^{p,p}(\mathbb{R}^{n})=\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$ is invariant under any Fourier integral operator of order zero as in Definition 2.12, if the symbol has compact support in the spatial variable. In this section we show that, by contrast, for all $q\in[1,\infty)\setminus\{p\}$, the space $\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n})$ is in general not invariant under such Fourier integral operators.
@@
For our example, let $\psi:\mathbb{R}^{2}\to\mathbb{R}^{2}$ be the diffeomorphism given by
@@
Any Fourier integral operator maps a wave packet, “located” at $(x,\omega)\in S^{*}\mathbb{R}^{n}$, to another (slightly distorted) wave packet, located at $\chi(x,\omega)$, where $\chi$ is the map on $S^{*}\mathbb{R}^{n}$ induced by the (homogeneous) canonical relation of the Fourier integral operator. Below, we will choose a function $u$ which is a sum of many wave packets all pointing in the same direction $\omega_{0}\in S^{1}$. We shall see that the $\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{2})$ norm of $u$ is then an $\ell^{p}$ sum of the $L^{p}(\mathbb{R}^{2})$ norms of the individual wave packets. Thus the growth of the norm in the number of wave packets $N$ will be approximately $N^{1/p}$. On the other hand, we have chosen $T$ such that the image of these wave packets will all have different directions. As a result, the $\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{2})$ norm of $Tu$ is an $\ell^{q}$ sum of the $L^{p}(\mathbb{R}^{2})$ norms of each of the translated wave packets, and as such the growth of the norm in $N$ of $Tu$ will be approximately $N^{1/q}$. Letting $N\to\infty$, this suffices for $q<p$.
@@
In applying the stationary phase expansion to (5.6), we need to check that we have a valid expansion as $R\to\infty$, i.e. that the powers of $R$ tend to $\infty$- ∞, since the amplitude in (5.6) also depends on $R$. The error estimate in [22, Theorem 7.7.5] after $m$ terms is a multiple of
@@
and this in turn is bounded above by a multiple of $R^{-m(1-2\gamma)}$, given that a derivative of the amplitude costs at worst $R^{\gamma}$. Now, since $\gamma$ was chosen to be strictly less than $1/2$, the power of $R$ in the error estimate tends to $\infty$- ∞, so this is a valid expansion as $R\to\infty$. For exactly the same reason, each term in the expansion, except for the zeroth one, has an overall power of $R$ that is strictly decreasing.
We aim to prove a lower bound on $\|v\|_{\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{2})}$. We claim that it suffices to prove a lower bound after replacing $v$ by its leading term in the stationary phase expansion of (5.6). In fact, due to symbolic derivative estimates on $\rho_{1}$ and $\Psi_{0}$, successive terms in this expansion have a similar form, but with decreasing powers of $R$. Each such term can be bounded above in the $\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{2})$ norm in a similar way as for $u$ above, and contributes less, by a factor of a negative power of $R$, than the leading term. On the other hand, the estimate on the remainder after taking $m$ terms in the stationary phase expansion of (5.6) is bounded by a multiple of $R^{-m(1-2\gamma)}$ for $\xi$ in a compact set, using [22, equation (7.7.12)], or by a multiple of $R^{2-m(1-\gamma)}\langle\xi\rangle^{-m}$ when $\langle\xi\rangle$ is large, using the “non-stationary phase” estimate of [22, equation (7.7.1)] (noting that $|d_{x}f|\eqsim\langle\xi\rangle$ for large $\langle\xi\rangle$). These pointwise estimates imply a bound on the $\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{2})$ norm of the remainder that decreases faster than any given power of $R$, if $m$ is taken sufficiently large. It follows that indeed, if we want to obtain a lower bound on $\|v\|_{\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{2})}$, valid for sufficiently large $R$, then (up to a factor of, say, $1/2$) we only need to consider the leading term of the expansion.
@@
for all $x\in\mathbb{R}^{2}$. We notice that the integrand is real and nonnegative, and it is bounded from below by a multiple of $R^{1/4}$ (since $\varphi_{\omega}$ is a symbol of order $1/4$, cf. property (2) in Section 3.1) in a rectangle of size approximately $R\times R^{1/2}$ with long axis in direction $\omega$. It follows that there is a rectangle in $x$-space centered at $x_{j,N}$, of size approximately $R^{-1}\times R^{-1/2}$, where $|v_{j,N,\mathrm{ldg}}(x)|$ is bounded from below by a multiple of $R^{7/4}$.
@@ Conclusion
Comparing (5.9) to (5.5), we see that the powers of $R$ match, but the power of $N$ in (5.5) is $1/p$, while the power of $N$ in (5.9) is $1/q$. As $R\to\infty$, we have $N^{1/q}\gg N^{1/p}$ (recall that $N\eqsim R^{\alpha}$). Given that we assumed that $q<p$, the map $T$ is therefore not bounded on $\mathcal{L}_{W,s}^{s,p,q}(\mathbb{R}^{2})$. ∎
@@
The proof above shows that the same phenomenon will happen for any Fourier integral operator whose canonical relation maps points $(x_{j},\omega_{0})$ in phase space, with varying base point $x_{j}$ and fixed direction $\omega_{0}$, to points with varying direction. For a concrete example, consider the Fourier integral operator $T=e^{it\sqrt{L}}$, where $L$ is the self-adjoint operator
@@
showing that $x_{1}$ and $\omega$ move along circles $\{x_{1}^{2}+\omega^{2}=r\}$ at a speed depending only on the radius $r$ of the circle. It is thus easy to write down the exact solution and verify that for a fixed small positive time $t$, if $\omega(0)=0$ and $x_{1}(0)=c$ then $\omega(t)=\omega(t,c)$ varies with $c$. Therefore, we can implement a similar example as above, to conclude that the Fourier integral operator $e^{it\sqrt{L}}$ is not bounded on $\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{2})$ for $t\neq 0$ and $q\neq p$. Again, we omit the details.
@@ 6.1. Preliminary results
We also remind the reader that $\varphi_{\omega}$ is a symbol of order $(n-1)/4$, so the multiplier $\langle D\rangle^{-\frac{n-1}{4}}\varphi_{\omega}(D)$ in (2) is a symbol of order zero.
@@ 6.2. Sobolev embeddings
Let $f$ be a local Hardy $\mathcal{H}^{1}(\mathbb{R}^{n})$ atom, $L^{\infty}$ normalized and adapted to a (Euclidean) ball of radius $r$. By translation invariance we may assume that the ball is centered at 00. First assume $10<r<$ 1, which is the main case. Then $\|f\|_{L^{\infty}(\mathbb{R}^{n})}\leq r^{-n}$, and $\int_{\mathbb{R}^{n}}f(x)\mathrm{d}x=0$. We will show that
@@
This completes our proof of (6.6) when $10<r<$ 1.
@@
is bounded from $H^{1}(\mathbb{R}^{n})$ into $L^{2}(S^{n-1};H^{1}(\mathbb{R}^{n}))$, the bound growing at most polynomially in $\operatorname{Im}z$. This is because growth in $\operatorname{Im}z$ arises from differentiating the kernel $K_{j}$, and the argument only requires a finite number of such derivatives. Furthermore, $T_{z}$ is bounded from $L^{2}(\mathbb{R}^{n})$ to $L^{2}(S^{n-1};L^{2}(\mathbb{R}^{n}))$ when $\textrm{Re}\,z=0$, by Plancherel, and this bound is independent of the imaginary part of $z$. Also note that if $\frac{1}{p}=\frac{1-\theta}{1}-\frac{\theta}{2}$, for $\theta\in(0,1)$ and $21<p<$ 2, then we have
@@
by e.g. [3, Theorems 5.1.1 and 5.1.2]). Thus by Stein’s complex interpolation theorem, $T_{-s(p)}$ is bounded from $L^{p}(\mathbb{R}^{n})$ to $L^{2}(S^{n-1};L^{p}(\mathbb{R}^{n}))$ for $21<p<$ 2. In other words,
@@
for $21<p<$ 2. Together with the estimate $\|\rho(D)f\|_{L^{p}(\mathbb{R}^{n})}\lesssim\|f\|_{W^{s(p),p}(\mathbb{R}^{n})}$ we obtain the assertion of Theorem 6.3 (1) for $21<p<$ 2. ∎

### v3

@@ 4.4. Interpolation and duality
$\mathcal{L}_{W,s_{0}}^{q_{0},p_{0}}(\mathbb{R}^{n}),\mathcal{L}_{W,s_{1}}^{q_{1},p_{1}}(\mathbb{R}^{n})]_{\theta}=\mathcal{L}_{W,s}^{q,p}(\mathbb{R}^{n}),$
@@
It suffices to consider the case where $s_{0}=s_{1}=s=0$. It is a basic fact about interpolation of vector-valued function spaces (see [26, Theorem 2.2.6]) that
$L^{q_{0}}_{\omega}L^{p_{0}}_{x}L^{2}_{\sigma}(S^{*}_{+}\mathbb{R}^{n}),L^{q_{1}}_{\omega}L^{p_{1}}_{x}L^{2}_{\sigma}(S^{*}_{+}\mathbb{R}^{n})]_{\theta}=L^{q}_{\omega}L^{p}_{x}L^{2}_{\sigma}(S^{*}_{+}\mathbb{R}^{n}).$
Hence one can combine Theorem 3.4 with a result about interpolation of complemented subspaces (see [41, Theorem 1.17.1.1]) to conclude that
$WVL^{q_{0}}_{\omega}L^{p_{0}}_{x}L^{2}_{\sigma}(S^{*}_{+}\mathbb{R}^{n}),WVL^{q_{1}}_{\omega}L^{p_{1}}_{x}L^{2}_{\sigma}(S^{*}_{+}\mathbb{R}^{n})]_{\theta}=WVL^{q}_{\omega}L^{p}_{x}L^{2}_{\sigma}(S^{*}_{+}\mathbb{R}^{n}).$
@@
$W:\mathcal{L}_{W,0}^{\tilde{q},\tilde{p}}(\mathbb{R}^{n})\rightarrow WVL^{\tilde{q}}_{\omega}L^{\tilde{p}}_{x}L^{2}_{\sigma}(S^{*}_{+}\mathbb{R}^{n})$
@@
Conversely, let $l\in(\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n}))^{*}$. Then $l\circ V\in(L^{q}_{\omega}L^{p}_{x}L^{2}_{\sigma}(S^{*}_{+}\mathbb{R}^{n}))^{*}$, by Corollary 4.4. Moreover, by [26, Theorems 1.3.10 and 1.3.21],
@@ 5.1. Boundedness of the Euclidean half-wave group
We first prove that our function spaces are invariant under the action of the Euclidean half-wave propagators. The following theorem contains the first statement of Theorem 1.1. Recall that $\mathcal{L}(X)$ is the collection of bounded operators on a Banach space $X$.
@@
Let $\rho^{\prime}\in C^{\infty}_{c}(\mathbb{R}^{n})$ be such that $\rho^{\prime}\equiv 1$ on $\mathrm{supp}(\rho)$, where $\rho$ is as in (4.1). It then follows from straightforward kernel estimates (see also [9, Theorem 1.18]) that
@@
for an implicit constant independent of $\omega\in S^{n-1}$ and $t\in\mathbb{R}$. To this end, we will adapt an argument from [16, Section 9] to the setting of anisotropic Hardy spaces.
@@
Remark 5.2.
It follows from [32, Proposition 3.2] that, for all $p,q\in[1,\infty)$ and $s\in\mathbb{R}$, one has $N\geq 2s(p)$ in Theorem 5.1. More precisely, here one can use that, up to constants which depend on $p$, $q$ and $s$, the $\mathcal{L}_{W,s}^{q,p}(\mathbb{R}^{n})$ norm of a function $f\in\mathcal{S}(\mathbb{R}^{n})$ with $\mathrm{supp}(\widehat{f}\,)\subseteq\{\xi\in\mathbb{R}^{n}\mid|\xi|\leq 1\}$ is equivalent to the $L^{p}(\mathbb{R}^{n})$ norm of $f$, as can for example be seen from (4.1).
Moreover, [32, Proposition 3.1] is an analogue of Theorem 5.1 for the Besov spaces associated with $\mathcal{L}_{W,s}^{q,p}(\mathbb{R}^{n})$, and there one has $N=2s(p)$. In future work, the same bound will be established in Theorem 5.1, at least for $p=q$.
@@ 5.2. Unboundedness of Fourier integral operators when $p\neq q$
We recall from [20] (see also [29, Corollary 2.4] and Corollary 6.6 below) that $\mathcal{L}_{W,s}^{p,p}(\mathbb{R}^{n})=\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$ is invariant under any Fourier integral operator of order zero as in Definition 2.12, if the symbol has compact support in the spatial variable. In this section we show that, by contrast, for all $q\in[1,\infty)\setminus\{p\}$, the space $\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{n})$ is in general not invariant under such Fourier integral operators.
@@
For our example, let $\psi:\mathbb{R}^{2}\rightarrow\mathbb{R}^{2}$ be the diffeomorphism given by
@@
Any Fourier integral operator maps a wave packet, “located” at $(x,\omega)\in S^{*}\mathbb{R}^{n}$, to another (slightly distorted) wave packet, located at $\chi(x,\omega)$, where $\chi$ is the map on $S^{*}\mathbb{R}^{n}$ induced by the (homogeneous) canonical relation of the Fourier integral operator. Below, we will choose a function $u$ which is a sum of many wave packets all pointing in the same direction $\omega_{0}\in S^{1}$. We shall see that the $\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{2})$ norm of $u$ is then an $\ell^{p}$ sum of the $L^{p}(\mathbb{R}^{2})$ norms of the individual wave packets. Thus the growth of the norm in the number of wave packets $N$ will be approximately $N^{1/p}$. On the other hand, we have chosen $T$ such that the image of these wave packets will all have different directions. As a result, the $\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{2})$ norm of $Tu$ is an $\ell^{q}$ sum of the $L^{p}(\mathbb{R}^{2})$ norms of each of the translated wave packets, and as such the growth of the norm in $N$ of $Tu$ will be approximately $N^{1/q}$. Letting $N\rightarrow\infty$, this suffices for $q<p$.
@@
In applying the stationary phase expansion to (5.6), we need to check that we have a valid expansion as $R\rightarrow\infty$, i.e. that the powers of $R$ tend to -$\infty$, since the amplitude in (5.6) also depends on $R$. The error estimate in [24, Theorem 7.7.5] after $m$ terms is a multiple of
@@
and this in turn is bounded above by a multiple of $R^{-m(1-2\gamma)}$, given that a derivative of the amplitude costs at worst $R^{\gamma}$. Now, since $\gamma$ was chosen to be strictly less than $1/2$, the power of $R$ in the error estimate tends to -$\infty$, so this is a valid expansion as $R\rightarrow\infty$. For exactly the same reason, each term in the expansion, except for the zeroth one, has an overall power of $R$ that is strictly decreasing.
We aim to prove a lower bound on $\|v\|_{\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{2})}$. We claim that it suffices to prove a lower bound after replacing $v$ by its leading term in the stationary phase expansion of (5.6). In fact, due to symbolic derivative estimates on $\rho_{1}$ and $\Psi_{0}$, successive terms in this expansion have a similar form, but with decreasing powers of $R$. Each such term can be bounded above in the $\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{2})$ norm in a similar way as for $u$ above, and contributes less, by a factor of a negative power of $R$, than the leading term. On the other hand, the estimate on the remainder after taking $m$ terms in the stationary phase expansion of (5.6) is bounded by a multiple of $R^{-m(1-2\gamma)}$ for $\xi$ in a compact set, using [24, equation (7.7.12)], or by a multiple of $R^{2-m(1-\gamma)}\langle\xi\rangle^{-m}$ when $\langle\xi\rangle$ is large, using the “non-stationary phase” estimate of [24, equation (7.7.1)] (noting that $|d_{x}f|\eqsim\langle\xi\rangle$ for large $\langle\xi\rangle$). These pointwise estimates imply a bound on the $\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{2})$ norm of the remainder that decreases faster than any given power of $R$, if $m$ is taken sufficiently large. It follows that indeed, if we want to obtain a lower bound on $\|v\|_{\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{2})}$, valid for sufficiently large $R$, then (up to a factor of, say, $1/2$) we only need to consider the leading term of the expansion.
@@
for all $x\in\mathbb{R}^{2}$. We notice that the integrand is real and nonnegative, and it is bounded from below by a multiple of $R^{1/4}$ (since $\varphi_{\omega}$ is a symbol of order $1/4$, cf. property (2) in Section 3.1) in a rectangle of size approximately $R\times R^{1/2}$ with long axis in direction $\omega$. It follows that there is a rectangle in $x$-space centered at $x_{j,N}$, of size approximately $R^{-1}\times R^{-1/2}$, where $|\varphi_{\omega}(D)v_{j,N,\mathrm{ldg}}(x)|$ is bounded from below by a multiple of $R^{7/4}$.
@@ Conclusion
Comparing (5.9) to (5.5), we see that the powers of $R$ match, but the power of $N$ in (5.5) is $1/p$, while the power of $N$ in (5.9) is $1/q$. As $R\rightarrow\infty$, we have $N^{1/q}\gg N^{1/p}$ (recall that $N\eqsim R^{\alpha}$). Given that we assumed that $q<p$, the map $T$ is therefore not bounded on $\mathcal{L}_{W,s}^{s,p,q}(\mathbb{R}^{2})$. ∎
@@
The proof above can be tweaked to show that the same phenomenon will occur for any Fourier integral operator with a canonical relation which maps points $(x_{j},\omega_{0})$ in phase space, with varying base point $x_{j}$ and fixed direction $\omega_{0}$, to points with varying direction. For a concrete example, consider the Fourier integral operator $T=e^{it\sqrt{L}}$, where $L$ is the self-adjoint operator
@@
showing that $x_{1}$ and $\omega$ move along circles $\{x_{1}^{2}+\omega^{2}=r\}$ at a speed depending only on the radius $r$ of the circle. It is thus easy to write down the exact solution and verify that for a fixed small positive time $t$, if $\omega(0)=0$ and $x_{1}(0)=c$ then $\omega(t)=\omega(t,c)$ varies with $c$. Therefore, we can implement a similar example as above, to conclude that the Fourier integral operator $e^{it\sqrt{L}}$ is not bounded on $\mathcal{L}_{W,0}^{q,p}(\mathbb{R}^{2})$ for $t\neq 0$ and $q\neq p$.
On the other hand, the positive result of Theorem 5.1 extends to FIOs $T$ that are Fourier multipliers. More precisely, if the phase function $\Phi$ is of the form
(5.10) $\Phi(x,\eta)=\lambda x\cdot\eta+\phi(\eta)$
for some $\lambda\in\mathbb{R}\setminus\{0\}$, a $\phi:\mathbb{R}^{n}\setminus\{0\}\rightarrow\mathbb{R}$ positively homogeneous of degree $1$, and all $(x,\eta)\in\mathbb{R}^{2n}\setminus o$, then $T$ maps wave packets in a fixed direction $\omega_{0}$ to wave packets in a (new) fixed direction $\omega_{1}$, and the conclusion of Theorem 5.1 still holds. This is analogous to [8, Theorem 5.2] which contains a similar property for the modulation spaces $M^{p,q}$. The close relation between the two results can be seen by observing that the conditions on the phase function in [8, Theorem 5.2], namely
$\sup_{x,x^{\prime},\eta}\big|\nabla_{x}\Phi(x,\eta)-\nabla_{x}\Phi(x^{\prime},\eta)\big|<\infty,$
together with our assumption that the phase function is homogeneous of degree $1$ in $\eta$, imply that $\Phi$ is an affine function (linear plus constant) in $x$, similar to (5.10). (On the other hand, the conditions imposed on the phase function in [8] are different from our conditions, and the authors of [8] do not require homogeneity of degree $1$ in $\eta$; indeed, as can be perceived from the other assumptions on the phase in [8, Theorem 5.2], their primary interest is the case of quadratic growth in $\eta$).
The same article, [8], also shows that a phenomenon analogous to Theorem 5.3 takes place on the modulation spaces $M^{p,q}$, in the sense that [8, Proposition 7.1] gives an example of an oscillatory operator which is bounded on $M^{p,p}$ but not on $M^{p,q}$ for any $p\neq q$. Their example is a multiplication operator with a symbol which has a quadratic (in $x$) oscillatory phase, and which is therefore not a Fourier integral operator in the sense of [23, 25] and of this article. The differences in the conditions on the phase function $\Phi$ between our article and [8] reflect the fact that their spaces and operators are designed for Schrödinger-like operators, whereas ours are designed for (half)-wave-like operators. In particular, the modulation spaces, apart from being more like Besov spaces than Hardy spaces, involve frequency localization to regions of unit size (reflecting the fact that the multiplier $|\eta|^{2}$ can only be linearized accurately on unit-size regions of frequency space), whereas the spaces $\mathcal{L}_{W,s}^{q,p}(\mathbb{R}^{n})$ localize at frequency level $R$ to dyadic-parabolic regions of dimension $R^{1/2}\times\ldots\times R^{1/2}\times R$ (reflecting the fact that the multiplier $|\eta|$ can be linearized accurately on such regions). This makes the analysis of the spaces, and the proofs of Theorem 5.3 and [8, Proposition 7.1], quite different.
@@ 6.1. Preliminary results
We also remind the reader that $\varphi_{\omega}$ is a symbol of order (n1)/4(n-1)/4, so the multiplier $\langle D\rangle^{-\frac{n-1}{4}}\varphi_{\omega}(D)$ in (2) is a symbol of order zero.
@@ 6.2. Sobolev embeddings
Let $f$ be a local Hardy $\mathcal{H}^{1}(\mathbb{R}^{n})$ atom, $L^{\infty}$ normalized and adapted to a (Euclidean) ball of radius $r$. By translation invariance we may assume that the ball is centered at 0. First assume $0<r<1$, which is the main case. Then $\|f\|_{L^{\infty}(\mathbb{R}^{n})}\leq r^{-n}$, and $\int_{\mathbb{R}^{n}}f(x)\mathrm{d}x=0$. We will show that
@@
This completes our proof of (6.6) when $0<r<1$.
@@
is bounded from $H^{1}(\mathbb{R}^{n})$ into $L^{2}(S^{n-1};H^{1}(\mathbb{R}^{n}))$, the bound growing at most polynomially in $\operatorname{Im}z$. This is because growth in $\operatorname{Im}z$ arises from differentiating the kernel $K_{j}$, and the argument only requires a finite number of such derivatives. Furthermore, $T_{z}$ is bounded from $L^{2}(\mathbb{R}^{n})$ to $L^{2}(S^{n-1};L^{2}(\mathbb{R}^{n}))$ when $\textrm{Re}\,z=0$, by Plancherel, and this bound is independent of the imaginary part of $z$. Also note that if $\frac{1}{p}=\frac{1-\theta}{1}+\frac{\theta}{2}$, for $\theta\in(0,1)$ and $1<p<2$, then we have
@@
by e.g. [3, Theorems 5.1.1 and 5.1.2]). Thus by Stein’s complex interpolation theorem, $T_{-s(p)}$ is bounded from $L^{p}(\mathbb{R}^{n})$ to $L^{2}(S^{n-1};L^{p}(\mathbb{R}^{n}))$ for $1<p<2$. In other words,
@@
for $1<p<2$. Together with the estimate $\|\rho(D)f\|_{L^{p}(\mathbb{R}^{n})}\lesssim\|f\|_{W^{s(p),p}(\mathbb{R}^{n})}$ we obtain the assertion of Theorem 6.3 (1) for $1<p<2$. ∎